What is the midpoint between (-1, 11) and (-1, -5)?
Answer by jim_thompson5910(35256)


Answer: (-1, 3)

==================================

Work Shown:

Add up the x coordinates to get
x1+x2 = -1+(-1) = -2
Then we cut this result in half
(-2)*(1/2) = -2/2 = -1
The x coordinate of the midpoint is x = -1 which is the same as the other original x coordinates.


Repeat similar steps for the y coordinates
Add: y1+y2 = 11+(-5) = 6
Divide by two: 6/2 = 3
The y coordinate of the midpoint is y = 3

Overall the midpoint is (-1, 3) which is shown by this diagram

We see that the distance from A = (-1,11) to B = (-1,-5) is exactly 11-(-5) = 11+5 = 16 units. Half of this is 16/2 = 8, which is the distance from A to C, or from B to C.

The midpoint C cuts segment AB in half, or bisects it. We can say AC = CB since both pieces are 8 units long.

Because the x coordinates of the original points are the same, we are effectively only concerned with finding the midpoint of the y coordinates. Imagine laying out a horizontal number line with -5 and 11 on it. Finding the midpoint should lead to 3. The number line can be placed vertically if you want, which is basically what the y axis visually represents.